OpenStudy (anonymous):

The area of a triangle whose height is 1 more than 6 times its base is 13 square feet. If the base of the triangle is x feet, which equation models this situation? A. 6x^2+x-12=0 B. 6x^2+x-13=0 C. 6x^2+x-26=0 D. 36x^2+6x-26=0

OpenStudy (anonymous):

@baru

OpenStudy (baru):

u know formula for area of a triangle?

OpenStudy (anonymous):

A=h*b

OpenStudy (baru):

\[ area=\frac{1}{2}bh\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

whats next

OpenStudy (baru):

ok, the question says area is 13 and that base is 'x' can you plug this information into the formula for area?

OpenStudy (anonymous):

a=13x?

OpenStudy (anonymous):

b?

OpenStudy (baru):

no

OpenStudy (anonymous):

a=1/2(13+x) ?

OpenStudy (baru):

area=13 base=b=x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it's not b?

OpenStudy (baru):

nah man...i cant just give you the final answer ;)

OpenStudy (anonymous):

x the Base. h is (6x+1) A =(1/2)bh = (1/2)x(6x+1) .......

OpenStudy (anonymous):

(1/2)x(6x+1) = 13 ft^2 x(6x+1) = 26 6x^2 + x - 26 = 0 (tossing out negative solution for unit measure) (6x + 13)(x - 2)= 0 x = 2ft, the Base. and h = 13ft

OpenStudy (anonymous):

And...checking (1/2)2ft•13ft = 13ft^2